# Theory and Atomistic Simulation of Electrodeposition

### v1

@@ 2.4 Electronucleation
When it comes to experimental electrochemical measurements, techniques such as chronoamperometry (which is a method where the current decay over time can provide information regarding the rate of cationic reduction and deposition onto an electrode surface) provide mostly macroscopic information, from which nanoscopic behaviour such as nucleation rates can be inferred 26, 112, 113. Such inferences, however, have been found to be inappropriate to describe the initial stages of nucleation where individual atoms and few-atom clusters are present 26, 25, 114, 115, 116, 117, 118 due to the assumptions within classical nucleation theory. Classical nucleation theory assumes the holding of the capillarity approximation, which treats the nucleus interior as a bulk incompressible liquid and assumes that the surface tension of a small liquid droplet is equal to the surface tension of a flat surface 119, 110, 111. This has been shown to break down for small systems 120, 121, 122. Furthermore, clusters are assumed to either grow or shrink via single-atom absorption or emission respectively, which places kinetic restrictions on the nucleation pathways 110, 111, 119. This does not hold in reality as entire clusters can merge or fragment, and these kinetic pathways cannot be ignored. While improvements to classical nucleation theory do exist, such as dynamical nucleation theory 123, mean-field kinetic nucleation theory 124, coupled flux theory 125, 126, 127, 128 and diffuse interface nucleation theory 129, 130, these have mostly been applied to describing the condensation of supersaturated vapors into the liquid phase and crystal nucleation studies rather than investigating metal electronucleation 119, 131. Despite its shortcomings, classical nucleation theory is still a powerful theory and has been shown to be capable of qualitatively capturing nucleation thermodynamics and kinetics for many systems 119.
The growth of metal nanostructures can be controlled either by kinetics, diffusion, or a combination thereof. Diffusion control typically occurs when there is a low concentration of metal cations within aqueous solutions 132. The growth of a new-phase hemispherical metal nanostructure of radius , under mixed kinetic-diffusion control, can be expressed using Equation (12). It should be noted that Equation (12) assumes that nucleation proceeds at moderate supersaturation levels and the aforementioned assumptions within classical nucleation theory are valid.
where are the current density and the exchange current density at the electrolyte/cluster interface respectively; are transfer coefficients such that ; is the overpotential; is the bulk concentration of depositing ions; and , where is the elementary electric charge 132. For purely kinetic-controlled growth, Equation (12) can be rearranged to form the Butler-Volmer equation 133, 132, which can be seen in Equation (13); it can also be rearranged to form an equation for pure diffusion-controlled growth in terms of the diffusion coefficient, as can be seen in Equation (14).
@@
In contrast, atomistic nucleation theories can be applied to clusters so small that is no longer continuous, as is the case with first-order phase transitions at high supersaturation () levels resulting in being non-differentiable 110, 111. Figure 2 shows the intrinsic differences between classical and atomistic theories of nucleation, as well as the (dis)continuity of cluster size. Atomistic theories allow for high levels to be modeled and have been validated against experimental studies 111. But despite the existence of such theories, much remains unclear regarding the initial stages of electronucleation and the role of the atomic-scale structure of the electrode surface. In this regard, explicit atomistic simulations can play an important role in elucidating the initial processes and mechanisms (vide infra).
3 Simulation of Electrochemical Reaction Conditions
@@
In this section, we summarize the various methodologies employed to simulate electrochemical reaction conditions. This includes an accurate account of the electrode potential, the electrolyte, and the electrode surface.
3.1 The Electrode Potential
@@
The electrode potential can greatly affect the reaction thermodynamics and kinetics during electrodeposition, and it thus becomes essential to accurately model the potential during atomistic simulations. Electrochemical experiments are typically performed under a constant potential and are referenced against well-defined reference electrodes. However, modeling a constant chemical potential within atomistic simulations is quite challenging, and various schemes have been developed to control the applied potential 134, 135, 48. The current schemes to model the applied electrode potential can be roughly split into three categories 136: classical forcefields 137, finite-field methods 138, 139, 140, 141 which can be used alongside forcefields or density functional theory (DFT), and grand-canonical ensembles (GCEs) with an electronic structure method 142, 143 such as DFT 144, 145. The first two methods require the full cell, while forcefields, in particular, do not treat electrons and can only treat electrostatics. Grand-canonical treatments, however, only require a half-cell to be modelled and can vary the number of electrons. The difference between full- and half-cell approaches is illustrated in Figure 3.
@@
Classical forcefield and finite-field approaches typically describe the applied electrode potential in terms of the inner potential difference between two electrodes, and do not take into account the electronic structure 146, 147, 148, 138, 139, 149, 150, 151, 137, 140, 141, 136. In both of these approaches, two electrode interfaces need to be simulated in order to build an electrode potential difference within the cell to enforce charge neutrality. However, as the properties of only the working electrode are of interest, the second electrode acts as a passive counter electrode. While the treatment of the additional electrode can be justified for forcefield potentials, it becomes computationally intractable for electronic structure methods, which computationally scale more severely than classical forcefields 152 with the number of atoms.
Forcefields come in a variety of flavors, including non-reactive and reactive, and often possess a trade-off between accuracy, via highly-detailed parameterizations, and transferability, ie. the number of materials to which they apply. Forcefields are widely used to simulate the energetics and dynamics of large systems by describing their nuclear interactions with a set of empirical equations rather than allowing for the electronic degrees of freedom that quantum mechanical methods explicitly specify. Recently, machine learning (ML) has emerged as a method to parameterize classical forcefields by supplying training data from quantum mechanical methods. Most ML-generated forcefields only account for local interactions though, which will not suffice for electrodeposition simulations, where long-range interactions have an important part to play, unless they are cleverly integrated alongside long-range interactions 153, 154, 155, 156, 157, 158, 159, 160, 161.
@@
In contrast, GCEs only require one electrode to be modeled and thus enable half-cells to be simulated. This is achieved by fixing the Fermi level (the energy level at which there is a 50% probability of it being occupied by an electron) of the electrode (which is equal to the chemical potential of electrons, ), while the chemical potential of the electrolyte, , is dependent on the electrolyte solution and its concentration 32. In order to have a well-defined treatment of an electrochemical solid-liquid interface, the electronic structure method (typically DFT) needs to be part of a GCE with a fluctuating number of electrons and ions (at a given temperature), as depicted in Figure 3, rather than the more common canonical ensemble, where the number of particles is constant but the chemical potentials are allowed to fluctuate 32. Practically, this is most elegantly achieved by fixing the Fermi level and allowing the number of electrons to fluctuate during a calculation 32. While stable algorithms for this method have been developed 31, 162, the fluctuating number of electrons can cause difficulties with convergence in calculations 32. An alternative methodology is to perform calculations at several points with a constant number of electrons and then interpolating to the desired Fermi level 163, 164, 165, 166, 167, 168, 169, 170.
@@
Changes in the Fermi level directly correspond to changes in , which is obtained by changing the charge state of the electrode. This presents a problem for simulations as electrochemical systems are usually partially periodic, but such systems need to be charge neutral. Various methodologies have been proposed to address this, including the introduction of a homogeneous background charge, correction schemes 164, 171, 172, joint DFT 173, 168, 31, and modified Poisson-Boltzmann implicit solvation models 174, 175, 176 (vide infra).
@@
An important quantity that is often used to analyze how changes in the electrode potential impact the electrochemical system is the electrostatic potential. The redistribution of electrons and ions in the system needed to maintain charge neutrality after changes in the electrode potential is reflected by changes to the electrostatic potential profile, particularly in the double layer region near the electrode. Figure 4 shows a typical electrostatic potential profile; at the cathode surface, the electrostatic potential typically exhibits oscillatory behavior due to the periodic arrangement of atoms and the resulting alternating regions of positive and negative charge density. In the bulk solution, the electrostatic potential tends to zero as the overall charge distribution is around neutral due to the lack of nearby charged surfaces and the cancellation of potentials from solvated ions. In the double layer, due to the presence of counterions close to the cathode, there is a sharp drop in the electrostatic potential often referred to as the Stern potential, which typically exhibits a large gradient near the cathode surface and decays towards zero as the potential extends towards the bulk solvent.
@@
Fermi level-fixed GC-DFT has been widely used to model electrochemical thermodynamics and kinetics. However, this is not suitable for outer-sphere reactions, semiconductors or two-electrode systems as in these systems, the Fermi level typically lies within the band gap rather than within the conduction or valence bands. To address the shortcomings of Fermi level-fixed GC-DFT, constant inner potential DFT has recently been proposed, which utilizes the local electrode inner potential as the thermodynamic parameter for the electrode potential, rather than the global Fermi level 136. Both GC-DFT variants have been shown to provide identical results for metallic electrodes 136, but differences can arise for semiconducting metal oxide–water interfaces 177.
3.2 The Electrolyte
@@
Before solvated cations are electrodeposited onto an electrode surface, they must diffuse through the solvent. Atomistic simulations must therefore ensure they appropriately describe the electrolyte, which includes the solvent itself and charged ions that, upon application of an electrochemical potential, form an electrochemical double-layer that modifies the electrostatic potential above the electrode surface 32.
@@
Figure 5 shows the various ways the solvent environment can be described for an electrode surface. Within atomistic simulations, the solvent environment can be described either by using explicit molecules, an implicit (analytical) model, or a hybrid of the two. In the former case, the explicit modeling of solvent molecules is a more ‘realistic’ and physically meaningful description of the system. However, explicit solvation models are typically more computationally expensive than implicit models and in order to model a physically meaningful system, many explicit solvent molecules need to be included within the model, and these can contribute to over 90% of the atoms within a modeled system 178. Using quantum mechanical methods to model solute-solute, solvent-solvent, and solute-solvent interactions can therefore quickly become computationally intractable with explicit solvent models comprising many hundreds of atoms. In order to reduce computational costs, empirical molecular mechanical forcefields have become a popular choice to treat interactions within the system. However, care must be taken during the parameterization of such forcefields in order to not sacrifice the accuracy that comes with ab initio approaches for the sake of computational tractability. In this regard, machine-learned interatomic potentials (MLIPs) offer a lot of promise 179, 180, assuming the existence of appropriate training data (vide infra).
@@
In most implicit solvent models, the solvent is treated as an electrostatic continuum with predefined dielectric and interfacial properties 181, 182, 183, 178, and the solute is placed inside a cavity within the continuum. In some cases, a dependency on distance from the solute can be included, and it is also possible to introduce a dependence on the rate of a particular process, whereby the response of the solvent varies for fast and slow processes. When using implicit solvent models, a number of choices need to be made, including the shape and size of the cavity that the solute will be placed within. Examples of simple cavity shapes are spherical and ellipsoidal, while more complex ones that can be generated algorithmically such as van der Waals surfaces (based on the van der Waals radii 184 of atoms), Lee-Richards molecular surfaces 185, Connolly surfaces 186, 187, and surfaces based on charge density isosurfaces 188 are also possible cavities.
@@
One of the advantages of implicit solvent models is that the number of interacting particles and the number of degrees of freedom within a system are significantly reduced. Implicit solvation models are thus typically computationally cheaper than explicit models and are therefore a good practical choice for computationally demanding studies 178, 189, 190. The reduced computational cost also means quantum mechanical methods become more tractable and can be used to treat the solute more accurately than typical molecular mechanical methods. However, the lack of an explicit atomistic description of the solvent can result in numerous interactions, such as hydrogen bonds (both with the solvent and within the solute), being neglected, overstabilized salt bridges, incorrect ion distribution 191, and unphysical sampling 178, 192, 193. Furthermore, the implicit description of the solvent also introduces an artificial boundary between the solute and solvent, and this interface needs to be treated carefully.
@@
Typically, the formation of the electrochemical double-layer is modeled using modified Poisson-Boltzmann implicit solvation models, which present the simplest level of fixed-potential GC-DFT for solid-liquid interfaces. Such models come in many varieties, including: linear, non-linear, and ion-size effects. With modified Poisson-Boltzmann solvers, charge neutrality can be maintained via the ionic distribution in the double-layer, though this constraint is not automatically fulfilled in non-linear models 194 due to an over-simplification of equations and the presence of the cavity exclusion function 32. This means that when combining modified Poisson-Boltzmann solvers with a quantum mechanical method for charged periodic systems, charge neutrality must be enforced either using Lagrange multipliers 194, a solvated jellium with a constant-charge background tempered by an implicit solvent 195, or by treating these systems using ‘metallic’ boundary conditions 163. With a solvated jellium model, a constant Fermi level is iteratively obtained and only selectively altered with respect to the eigen-levels of the adsorbed solvent 195.
@@
Finally, hybrid solvation models that seek to reap the benefits of both explicit and implicit solvation models can also be used to effectively model the electrolyte. In terms of the modeling, hybrid models can be facilitated using embedded cluster models (vida infra) and methodologies such as hybrid quantum mechanics/molecular mechanics (QM/MM). Here, the system can be partitioned into three regions, where the central region is treated using QM and comprises the solute and some explicit solvent molecules. This central part is embedded within a second layer and contains more explicit solvent molecules but is treated using MM. Finally, both aforementioned regions are embedded within a third layer, where the solvent is described using implicit models and represents the bulk solvent. This hybrid approach allows the local region of interest to be modeled with the accuracy of QM without subjecting the entire system to the typically higher computational costs that come with QM. However, care must be taken to ensure the various interfaces between the regions are modeled appropriately, and some studies have also observed a dependency on the number of added explicit solvent molecules 196.
3.3 The Electrode Surface
3.3.1 The Electronic Structure Method
@@
Kohn-Sham DFT 144, 145 is one of the most commonly used electronic structure methods to describe extended surfaces 197 and materials 198. Within DFT, increasingly accurate density-functional approximations are being developed 199 that can represent the energetics and electronic structure of complex materials. To describe electrified solid-liquid interfaces efficiently, a pragmatic selection of well-tested density-functional approximations that balance computational efficiency and predictive accuracy is required 198, 197. For example, a posteriori long-range dispersion corrections, such as the Grimme 200, 201, 202, 203, 204 and Tkatchenko-Scheffler 205, 206, 207, 208, 209, 210, 211, 212 families, to generalized gradient approximations have been shown to provide a reliably accurate representation of adsorption structures and energetics. However, considering the existing limitations in the quantitative experimental characterization of electrochemical systems in general, and the kinetics and dynamics of electrodeposition in particular, the accuracy of existing density-functional approximations and dispersion correction schemes is currently not the limiting factor in atomistic electrodeposition simulations. Nevertheless, the choice of quantum mechanical method used remains key for the accurate modeling of processes at electrode surfaces.
@@
However, the intrinsic computational scalability of DFT and even beyond-DFT, such as wavefunction methods and many-body perturbation theory, provides a challenge for systems comprising more than a few hundred atoms, which can often be the case for complex electrode surfaces. DFT typically has a formal scaling, while higher-level ab initio wavefunction methods can formally scale beyond (where is the number of basis functions). Semi-empirical tight-binding methods, such as density functional tight-binding 213, have grown in popularity as they comprise a good compromise between computational cost and accuracy and have been successfully employed to study the electrodeposition of metals 214. However, tight-binding parameterizations are typically developed for a particular subset of elements for a specific purpose, which means they possess low transferability. Furthermore, only a few reliable parameterizations therefore currently exist for metal-organic interfaces 215, 216, 217, 214, which can be problematic for describing metal electrodeposition on carbonaceous electrodes, such as graphite/graphene and boron-doped diamond, or even for multimetallic systems. For these reasons, DFT is still typically more popular than tight-binding methods. However, newer semi-empirical methods, such as xTB, have sought to correct for this drawback by employing a global and element-specific parameterization 218, 219, rather than the element pair-specific parameters employed within density functional tight-binding.

### v2

@@ Introduction
When it comes to modeling the process of electronucleation, both classical and atomistic theories exist to describe the formation of stable nuclei 26, 126, 127, 128. Classical nucleation theory relies on macroscopic physical quantities that are applicable to sufficiently large clusters such that their size can be considered a continuous variable. Common experimental electrochemistry techniques typically provide mostly macroscopic information, from which nanoscopic behavior such as nucleation rates are inferred 26, 129, 130. Such inferences, however, have been found to be inappropriate to describe the initial stages of nucleation where individual atoms and few-atom clusters are present 26, 25, 131, 132, 133, 134, 135. Classical nucleation theory assumes that the capillarity approximation holds 136, 127, 128, which has been shown to break down for small systems 137, 138, 139. Furthermore, clusters are assumed to either grow or shrink via single-atom absorption or emission, respectively, which places kinetic restrictions on the nucleation pathways 127, 128, 136. This does not hold in reality as entire clusters can merge or fragment, and these kinetic pathways cannot be ignored. While improvements to classical nucleation theory do exist, such as dynamical nucleation theory 140, mean-field kinetic nucleation theory 141, coupled flux theory 142, 143, 144, 145 and diffuse interface nucleation theory 146, 147, these have mostly been applied to describing the condensation of supersaturated vapors into the liquid phase and crystal nucleation studies rather than investigations of metal electronucleation 136, 148. Despite its shortcomings, classical nucleation theory is still a powerful theory and has been shown to be capable of qualitatively capturing nucleation thermodynamics and kinetics for many systems 136.
@@
In contrast, atomistic nucleation theories can be applied to clusters so small that their size can no longer be considered to be continuous, as is the case with first-order phase transitions at high supersaturation levels 127, 128. Atomistic theories allow for high supersaturation levels to be modeled and have been validated against experimental studies 128. Despite the existence of such theories, much remains unclear regarding the initial stages of electronucleation and the role of the atomic-scale structure of the electrode surface. In this regard, explicit atomistic simulations can play an important role in elucidating the initial processes and mechanisms (vide infra).
The electronucleation of metallic phases is inherently a multiscale problem, spanning atomic-scale processes to macroscopic current transients. A well-established approach to describing electronucleation kinetics involves the analysis of current transients at constant potential, which can provide insights into electronucleation rates and cluster growth dynamics. This formulation, developed through the seminal works of Scharifker 149, 150, 151, 152 and Milchev 129, 128, 127, has provided crucial theoretical frameworks for understanding electronucleation mechanisms. Their models distinguish between instantaneous and progressive electronucleation, describing how nuclei emerge and evolve over time based on the electrochemical conditions. While these approaches have been reviewed previously 153, revisiting these key developments is important for contextualizing more recent advancements in atomistic modeling and experimental validation. Linking macroscopic observables, such as current transients, with atomistic descriptions remains a central challenge, highlighting the necessity of multiscale modeling approaches to bridge theoretical and experimental perspectives on electronucleation.
Open Questions Driving the Need for New Simulation Methods
Despite significant advances in both experimental and computational approaches, a complete and predictive understanding of electrodeposition remains elusive. A major challenge is the current disconnect between atomistic and macroscopic models, which limits the ability to describe the entire deposition process, from the initial stages of adsorption and charge transfer to mesoscopic structure formation and macroscopic film growth. As discussed below, atomistic simulations can provide detailed insights into local interactions, electron transfer 118, and electronucleation at the atomic scale, but are often computationally demanding and struggle to capture long timescale dynamics and large system sizes relevant to real electrodeposition processes. Meanwhile, continuum models 154, 155, 156, 157, 158 (vide infra) can simulate extended growth patterns but often rely on parameters that lack direct links to the underlying atomic-scale phenomena. Bridging this scale gap remains a critical open problem.
Furthermore, current simulation methods need to incorporate realistic electrochemical environments, such as the influence of solvent dynamics, and the structure of the double layer. Another key challenge is capturing on-surface electronucleation and growth mechanisms with sufficient accuracy. Explicit modeling of the formation of stable growth nuclei and their kinetics remains an area where more computational activity is needed. In addition, the morphology of the electrode can include numerous complexities, such as nanoscale heterogeneities, surface defects, and grain boundaries, which all play a role in localizing charge transfer and influencing electronucleation kinetics. Experimental studies have shown that surface morphology and defect concentration can dramatically alter Tafel slopes, suggesting a direct link between electronucleation events on the atomic scale and macroscopic electrochemical behavior. However, current theoretical models often assume idealized surfaces, limiting their ability to predict realistic deposition kinetics. Developing atomistic models that explicitly account for surface roughness and defects and their impact on charge transfer kinetics remains an important challenge.
A key challenge in understanding electrodeposition lies in the disconnect between kinetic theory and experimentally observed kinetics. Kinetic models, such as Butler-Volmer and Marcus-Hush-Chidsey formalisms, describe charge transfer kinetics under idealized assumptions, often neglecting critical factors such as ion solvation, surface heterogeneities, and mass transport effects. Experimentally measured Tafel slopes often deviate from theoretical predictions due to these complexities, particularly under operando conditions where solvent interactions and dynamic restructuring of the electrochemical interface play critical roles. Ab initio atomistic simulation methods offer a powerful bridge between fundamental kinetic theory and experimental observables and would enable the direct computation of reaction barriers, overpotential-dependent activation energies, and coupled ion-electron transfer mechanisms. By parameterizing kinetic models with ab initio data, a more accurate description of reaction pathways and transfer coefficients can be achieved. For example, ab initio simulations could quantify how solvation and desolvation contribute to the potential energy landscape, refining the description of coupled ion-electron transfer theory beyond empirical parameterization. Moreover, ab initio calculations could provide insights into the potential-dependent nature of the transfer coefficient by explicitly modeling the influence of electronic structure and reorganization energy at the electrode interface. Such an approach would permit a direct comparison with experimental Tafel slopes and voltammetric data, potentially resolving long-standing discrepancies between theoretical predictions and observed electrochemical kinetics. A multiscale approach, via integration of atomistic simulations with experimental measurements, could therefore play a crucial role in advancing understanding of electrodeposition processes and improving the predictive power of kinetic models.
Developing better simulation methods will not only improve computational predictions but also provide valuable guidance for experimentalists. Simulations can help interpret complex experimental observations and predict electrodeposition conditions that lead to desirable morphologies or the suppression of defects. By integrating atomistic insights with macroscale modeling, new computational approaches have the potential to unlock more precise control over electrodeposition processes, ultimately advancing technologies in areas such as electrode manufacturing, nanostructured catalysts, and corrosion-resistant coatings. Addressing these open questions will require a concerted effort in developing multiscale modeling frameworks, improved parameterization strategies, and greater synergy between simulations and experiments.
Simulation of Electrochemical Reaction Conditions
@@
In this section, we summarize the various state-of-the-art methodologies employed to simulate electrochemical reaction conditions. This includes an accurate account of the electrode potential, the electrolyte, and the electrode surface.
The Electrode Potential
@@
The electrode potential can greatly affect the reaction thermodynamics and kinetics during electrodeposition, and it thus becomes essential to accurately model the potential during atomistic simulations. Electrochemical experiments are typically performed under a constant potential and are referenced against well-defined reference electrodes. However, modeling a constant chemical potential within atomistic simulations is challenging, and various schemes have been developed to control the applied potential 159, 160, 76. The current schemes to model the applied electrode potential can be roughly split into three categories 161: classical force fields 162, finite-field methods 163, 164, 165, 166 which can be used alongside force fields or density functional theory (DFT), and grand-canonical ensembles (GCEs) with an electronic structure method 167, 168 such as DFT 169, 170. The first two methods require the full cell. Force fields do not treat electrons and can only treat electrostatic effects. Grand-canonical treatments, however, only require a half-cell to be modeled with the number of electrons controlled by the applied electrochemical potential. The difference between full- and half-cell approaches is illustrated in Figure 2.
@@
Classical force field and finite-field approaches typically describe the applied electrode potential in terms of the inner potential difference between two electrodes, and do not take into account the electronic structure 171, 172, 173, 163, 164, 174, 175, 176, 162, 165, 166, 161. In both of these approaches, two electrode interfaces need to be simulated in order to build an electrode potential difference within the cell to enforce charge neutrality. However, as the properties of only the working electrode are of interest, the second electrode acts as a passive counter electrode. While the treatment of the additional electrode can be justified for molecular dynamics simulations with force fields, it becomes computationally intractable for electronic structure methods, which do not scale as favorably in terms of their computational cost with the number of atoms in the system 177. Force fields come in a variety of flavors, including non-reactive and reactive, and often possess a trade-off between accuracy, via highly-detailed parameterizations, and transferability, i.e. the number of materials to which they apply. Force fields are widely used to simulate the energetics and dynamics of large systems by describing their interatomic and intermolecular interactions with a set of parametrized analytical functions. Recently, machine learning (ML) methods have emerged as a successful approach to parameterize force fields by supplying training data from ab initio calculations. 178, 179
@@
In contrast, GCE simulations only require one electrode to be modeled and thus enable half-cells to be simulated. This is achieved by fixing the Fermi level of the electrode (which is equal to the chemical potential of electrons, ), while the chemical potential of the electrolyte, , is dependent on the electrolyte solution and its concentration 60. To achieve a well-defined treatment of an electrochemical solid-liquid interface, the electronic structure method (typically DFT) needs to be part of a GCE with a fluctuating number of electrons and ions (at a given temperature), as depicted in Figure 2, rather than the more common canonical ensemble, where the number of particles is constant but the chemical potentials are allowed to fluctuate 60. Practically, this is most elegantly achieved by fixing the Fermi level and allowing the number of electrons to fluctuate during a calculation 60. While stable algorithms for this method have been developed 59, 180, the fluctuating number of electrons can cause difficulties with convergence in calculations 60. An alternative methodology is to perform calculations with different fixed numbers of electrons and then interpolate to the desired Fermi level 181, 182, 183, 184, 185, 186, 187, 188.
@@
Changes in the Fermi level directly correspond to changes in , which is obtained by changing the charge state of the electrode. This presents a problem for simulations as electrochemical systems are usually partially periodic, but such systems need to be charge neutral. Various methodologies have been proposed to address this, including the introduction of a homogeneous background charge, correction schemes 182, 189, 190, joint DFT 191, 186, 59, and modified Poisson-Boltzmann implicit solvation models 192, 193, 194.
@@
An important quantity that is often used to analyze how changes in the electrode potential impact the electrochemical system is the electrostatic potential. The redistribution of electrons and ions in the system needed to maintain charge neutrality after changes in the electrode potential is reflected by changes to the electrostatic potential profile, particularly in the double layer region near the electrode. Figure 3 shows a simplified electrostatic potential profile; at the cathode surface, the electrostatic potential typically exhibits oscillatory behavior due to the periodic arrangement of atoms and the resulting alternating regions of positive and negative charge density. In the bulk solution, the electrostatic potential tends to zero as the overall charge distribution is around neutral due to the lack of nearby charged surfaces and the cancellation of potentials from solvated ions. However, beyond the metal surface, the electrostatic potential will typically exhibit damped oscillations that reflect the presence of distinct solvation shells around the metal ions. While it can be difficult to assign the exact location of the double layer given the broad continuous nature of the electrostatic potential 195, we provide a simplistic visual indication of the double layer region in Figure 3.
@@
Fermi level-fixed GC-DFT has been widely used to model electrochemical thermodynamics and kinetics. However, this is not suitable for outer-sphere reactions, semiconductors or two-electrode systems as in these systems, the Fermi level typically lies within the band gap rather than within the conduction or valence bands. To address the shortcomings of Fermi level-fixed GC-DFT, constant inner potential DFT has recently been proposed, which utilizes the local electrode inner potential as the thermodynamic parameter for the electrode potential, rather than the global Fermi level 161. Both GC-DFT variants have been shown to provide identical results for metallic electrodes 161, but differences can arise for semiconducting metal oxide–water interfaces 196.
The Electrolyte
@@
Before solvated cations are electrodeposited onto an electrode surface, they must diffuse through the solvent. Atomistic simulations must therefore ensure they appropriately describe the electrolyte, which includes the solvent itself and charged ions that, upon application of an electrochemical potential, form an electrochemical double-layer that modifies the electrostatic potential above the electrode surface 60.
@@
Figure 4 shows the various ways the solvent environment can be described for an electrode surface. Within atomistic simulations, the solvent environment can be described either by using explicit molecules, an implicit (analytical) model, or a hybrid of the two. In the former case, the explicit modeling of solvent molecules is a more ‘realistic’ and physically meaningful description of the system. Specifically in the context of electrodeposition, counter ions such as chloride anions will be present at high concentrations and may need to be explicitly treated. Close to the cathode they will strongly affect the electrostatic interactions, polarization and solvation dynamics during electrodeposition. However, explicit solvation models are typically more computationally expensive than implicit models and in order to model a physically meaningful system, many explicit solvent molecules need to be included within the model, and these can contribute to over 90% of the atoms within a modeled system 197. Using quantum mechanical methods to model solute-solute, solvent-solvent, and solute-solvent interactions can therefore quickly become computationally intractable with explicit solvent models comprising many hundreds of atoms. To reduce computational costs, empirical molecular mechanical force fields have become a popular choice to treat interactions within the system. However, care must be taken during the parameterization of such force fields in order to not sacrifice the accuracy that comes with ab initio approaches for the sake of computational tractability. In this regard, machine-learned interatomic potentials (MLIPs) offer a lot of promise 198, 199.
@@
In most implicit solvent models, the solvent is treated as an electrostatic continuum with predefined dielectric and interfacial properties 200, 201, 202, 197, and the solute is placed inside a cavity within the continuum. In some cases, a dependency on distance from the solute can be included, and it is also possible to introduce a dependence on the rate of a particular process, whereby the response of the solvent varies for fast and slow processes. When using implicit solvent models, a number of choices need to be made, including the shape and size of the cavity that the solute will be placed within. Examples of simple cavity shapes are spherical and ellipsoidal, while more complex ones that can be generated algorithmically such as van der Waals surfaces (based on the van der Waals radii 203 of atoms), Lee-Richards molecular surfaces 204, Connolly surfaces 205, 206, and surfaces based on charge density isosurfaces 207 are also possible cavities.
@@
One of the advantages of implicit solvent models is that the number of interacting particles and the number of degrees of freedom within a system are significantly reduced. Implicit solvation models are thus typically computationally cheaper than explicit models and are therefore a good practical choice for computationally demanding studies 197, 208, 209. The reduced computational cost also means quantum mechanical methods become more tractable and can be used to treat the solute more accurately than typical molecular mechanical methods. However, the lack of an explicit atomistic description of the solvent can result in numerous interactions, such as hydrogen bonds (both with the solvent and within the solute), being neglected, overstabilized salt bridges, incorrect ion distribution 210, and unphysical sampling 197, 211, 212. Furthermore, the implicit description of the solvent also introduces an artificial boundary between the solute and solvent, and this interface needs to be treated carefully.
@@
Typically, the formation of the electrochemical double-layer is modeled using modified Poisson-Boltzmann implicit solvation models, which present the simplest level of fixed-potential GC-DFT for solid-liquid interfaces. Such models come in many varieties, including: linear, non-linear, and with incorporation of ion-size effects. With modified Poisson-Boltzmann solvers, charge neutrality can be maintained via the ionic distribution in the double-layer, though this constraint is not automatically fulfilled in non-linear models 213 due to an over-simplification of equations and the presence of the cavity exclusion function 60. This means that when combining modified Poisson-Boltzmann solvers with a quantum mechanical method for charged periodic systems, charge neutrality must be enforced either using Lagrange multipliers 213, a solvated jellium with a constant-charge background tempered by an implicit solvent 214, or by treating these systems using ‘metallic’ boundary conditions 181. With a solvated jellium model, a counter charge is dispersed in implicit solvent and a fixed electrochemical potential is iteratively obtained by varying the number of electrons. This works well for metal electrodes, where excess electrons typically localize at the electrode surface and change the work function. 214.
@@
Finally, hybrid solvation models that seek to reap the benefits of both explicit and implicit solvation models can also be used to effectively model the electrolyte. In terms of the modeling, hybrid models can be facilitated using embedded cluster models (vida infra) and methodologies such as hybrid quantum mechanics/molecular mechanics (QM/MM). Here, the system can be partitioned into three regions, where the central region is treated using QM and comprises the solute and some explicit solvent molecules. This central part is embedded within a second layer and contains more explicit solvent molecules but is treated using MM. Finally, both aforementioned regions are embedded within a third layer, where the solvent is described using implicit models and represents the bulk solvent. This hybrid approach allows the local region of interest to be modeled with the accuracy of QM without subjecting the entire system to the typically higher computational costs that come with QM. However, care must be taken to ensure the various interfaces between the regions are modeled appropriately, and some studies have also observed a dependency on the number of added explicit solvent molecules 215.
The Electrode Surface
The Electronic Structure Method
@@
Kohn-Sham DFT 169, 170 is one of the most commonly used electronic structure methods to describe extended surfaces 216 and materials 217. Within DFT, increasingly accurate density-functional approximations are being developed 218 that can represent the energetics and electronic structure of complex materials. To describe electrified solid-liquid interfaces efficiently, a pragmatic selection of well-tested density-functional approximations that balance computational efficiency and predictive accuracy is required 217, 216. For example, a posteriori long-range dispersion corrections, such as the Grimme 219, 220, 221, 222, 223 and Tkatchenko-Scheffler 224, 225, 226, 227, 228, 229, 230, 231 families, to generalized gradient approximations have been shown to provide a reliably accurate representation of adsorption structures and energetics. However, considering the existing limitations in the quantitative experimental characterization of electrochemical systems in general, and the kinetics and dynamics of electrodeposition in particular, the accuracy of existing density-functional approximations and dispersion correction schemes is currently only one of several limiting factor in atomistic electrodeposition simulations. Nevertheless, the choice of quantum mechanical method used remains key for the accurate modeling of processes at electrode surfaces.
@@
The intrinsic computational scalability of DFT and even beyond-DFT, such as wavefunction methods and many-body perturbation theory, provides a challenge for systems comprising more than a few hundred atoms, which can often be the case for complex electrode surfaces. Semi-empirical tight-binding methods, such as density functional tight-binding (DFTB) 232, have grown in popularity as they comprise a good compromise between computational cost and accuracy and have been successfully employed to study the electrodeposition of metals 233. However, tight-binding parameterizations are typically developed for a particular subset of elements for a specific purpose, which means they provide low transferability across systems. Furthermore, only a few reliable parameterizations currently exist for metal-organic interfaces 234, 235, 236, 233, which can be problematic for describing metal electrodeposition on graphite/graphene and boron-doped diamond, or even on multimetallic systems. For these reasons, DFT is still typically more popular than tight-binding methods. However, newer semi-empirical methods, such as xTB, have sought to correct for this drawback by employing a global and element-specific parameterization 237, 238, rather than the element pair-specific parameters employed within DFTB.